OpenStudy (surryyy):

Which point is on the line of the equation y+3/4=-2(x+2) ? (-2,2) (2, 3/4) (-2, -3/4) (4, 3/4)

OpenStudy (sohailiftikhar):

(-2,-3/4)

OpenStudy (sohailiftikhar):

you can consider them as x1=-2 and y1=-3/4 now slope =y-y1/x-x1 y-(-3/4)/x-(-2)=-2 y+3/4=-2(x+2)

OpenStudy (surryyy):

thnx